# Why are Visually-Grounded Language Models Bad at Image Classification?

### v1

@@ Abstract
Image classification is one of the most fundamental capabilities of machine vision intelligence. In this work, we revisit the image classification task using visually-grounded language models (VLMs) such as GPT-4V and LLaVA. We find that existing proprietary and public VLMs, despite often using CLIP as a vision encoder and having many more parameters, significantly underperform CLIP on standard image classification benchmarks like ImageNet. To understand the reason, we explore several hypotheses concerning the inference algorithms, training objectives, and data processing in VLMs. Our analysis reveals that the primary cause is data-related: critical information for image classification is encoded in the VLM’s latent space but can only be effectively decoded with enough training data. Specifically, there is a strong correlation between the frequency of class exposure during VLM training and instruction-tuning and the VLM’s performance in those classes; when trained with sufficient data, VLMs can match the accuracy of state-of-the-art classification models. Based on these findings, we enhance a VLM by integrating classification-focused datasets into its training, and demonstrate that the enhanced classification performance of the VLM transfers to its general capabilities, resulting in an improvement of 11.8% on the newly collected ImageWikiQA dataset.
@@ 2 VLMs are Bad at Image Classification
2.1 Models
VLMs.
CLIPs.
2.2 Data
2.3 Evaluation Protocol
VLMs.
CLIPs.
2.4 Results
@@ 3 Why are VLMs Bad Image Classifiers?
    Finally, we investigate data-related questions. For example, does the VLM training data include enough classification data and cover enough classes? We find a strong correlation between class exposure in training and model performance. Moreover, VLMs can achieve the same level of performance as CLIPs when trained with enough data. These results suggest that data is the primary cause of the poor classification performance of VLMs.
@@ 3.1 Inference
Table 2: Analysis of VLMs from the inference perspective. (Top) We explore prompt variation such as wording, label order, chain-of-thought and find it has limited impact on the performance. (Bottom) We leverage the probabilistic inference strategy, which improves the performance but still fails to close the gap between VLMs and CLIPs.
@@ 3.2 Training
Table 3: Analysis of VLMs from the training perspective. (Left) We conduct feature probing experiments on the VLM’s last layer and find that the information required for classification is mostly preserved in the VLM’s latent space. (Right) We fine-tune VLMs on the classification datasets using the text generation objective and find that the text generation training objective is as effective as the traditional cross-entropy for learning classification, which eliminates VLM-CLIP performance gap.
@@
We find that the text generation training objective is as effective as the traditional cross-entropy for learning classification tasks, which eliminates the performance gap between VLMs and CLIPs. From Table 3, we find that LLaVA1.5-7B achieves 85.7% accuracy on ImageNet, same as 85.2% accuracy when fine-tuning its vision encoder CLIP-L. The same findings apply to all the VLMs on all the datasets.
@@ 3.3 Data
As observed, the visual information is preserved in the VLM’s latent space, and the VLM’s training objective is sufficient for learning classification tasks. We hypothesize that the poor classification performance of VLMs is due to their training data. For example, there might be insufficient classification data or a lack of diverse classes. To investigate this, we analyze the LLaVA1.5-7B training data [29], the only fully publicly available VLM training dataset. We examined the relationship between the frequency of class occurrence and the VLM’s classification performance on those classes***LLaVA training contains two stages. Here we combine the data from the two stages. More in Appendix B.6..
Our findings indicate that data determines VLM classification performance. As shown in Figure 3, there is a strong correlation between the presence of class labels in the VLM training data and the VLM’s classification accuracy for those classes. The LLaVA1.5-7B model achieves 82.7% zero-shot accuracy on ImageNet classes with more than 10,000 occurrences in its training data, but only 3.0% accuracy on classes with fewer than 10 occurrences. The Spearman correlation between class frequency and class performance is very high (0.76 on ImageNet; see Appendix §B.6). In contrast, there is no correlation between its vision encoder CLIP-L’s performance on those classes with the same VLM training data, and after fine-tuning LLaVA1.5-7B on the ImageNet classification data, the strong correlation disappears. Combining all of our results, we conclude that data plays a critical role in determining the VLM’s classification performance. In Appendix §D, we further show that the data type (e.g., classification or captioning data of a given class) is unimportant.
@@ 4.2 ImageNetQA
To bridge the gap between classification and advanced capabilities, we introduce ImageWikiQA, an object-centric, knowledge-intensive question answering dataset that combines both worlds. Each question in ImageWikiQA is a multiple-choice question with four options and one correct answer. Although each question does not directly ask for the category of the object within the image, the question can only be accurately answered if the class of the object is correctly identified. Example questions from ImageWikiQA can be found in Figure 1 and Appendix §C.1.
@@ Acknowledgements
Broader Impacts and Ethics Statement
In this work, we identify a core limitation of visually-grounded language models (VLMs) and propose a simple solution to address this issue. We find that VLMs underperform in image classification. By integrating classification data into VLM training, we have demonstrated that this intervention not only enhances classification accuracy but also improves the general capabilities of VLMs. This advancement holds promise for real-world applications, such as virtual assistants for visually impaired individuals. However, there are potential risks associated with incorporating race-based or gender-based classification data, which could amplify discriminatory biases in VLMs. We strongly emphasize the importance of ethical and responsible use of our approach to ensure it contributes positively to our digital ecosystem.
Reproducibility Statement
We provide an open-source implementation of our work and have released the ImageWikiQA dataset at https://github.com/yuhui-zh15/VLMClassifier. This will enable researchers to reproduce all experiments described in this paper and conduct their own analyses on additional VLMs and datasets.
Limitations
Due to computational constraints, we cannot evaluate all possible VLMs and datasets. We select two representative ones, LLaVA and BLIP, and conduct analyses on four datasets. Additionally, as we find that VLMs have information in their latent space, we conduct large-scale training to decode them. Developing a zero-shot method to decode this information without extensive training could be a promising direction for future research.
Compute Resources
We use four NVIDIA L40S GPUs for all experiments. Most inference experiments take 5-20 hours on a single L40S GPU. The most computationally intensive step is fine-tuning. Fine-tuning LLaVA1.5-7B on 1.28M ImageNet data takes 1.4 days for one epoch using four GPUs. Fine-tuning LLaVA1.5-7B on 1.28M ImageNet data combined with 665K original LLaVA instruction-tuning data takes 3.5 days for one epoch using four GPUs. For the 13B model, the cost is approximately 40% higher.
Summary of Appendix
We provide additional details for each section in the main paper.
  • In §A, we provide details of models and data for §2.
  • In §B, we provide details of experimental settings and results for §3.
  • In §C, we provide details of our collected ImageNetQA dataset for §4.
  • In §D, we discuss how data types affect VLM classification performance.
Appendix A Supplementary Section 2
@@ B.1 Prompt Variation Detail
For the fixed order, we use the default dataset order for each image.
@@ B.2 Label Set Size Details
The table version of Figure 2 is provided in Table 7.
@@
We observe that while the absolute performance gap between VLMs and CLIP decreases when reducing the number of classes, the relative performance gap increases. This trend is indicated by the error rates between VLMs and CLIP shown in Figure 4.
@@ C.2 Results Details
Appendix D Ablation on Data Types
In the main paper, we have demonstrated that data is crucial for classification performance for visually-grounded language models (VLMs) and proposed a simple data integration method to enhance VLMs. However, we are curious whether the data type also plays a significant role in VLM classification performance. For instance, to train a VLM to classify “guinea pig”, should the model be trained with classification-focused data (e.g., “<image> Question: What type of object is in this photo? Answer: Guinea pig.”), or can it be trained with any type of data, such as caption-focused data (e.g., “<image> Question: Generate a caption for the image. Answer: A guinea pig playing in the grass.”) or VQA-focused data (e.g., “<image> Question: What is the native region of this guinea pig? Answer: South America.”)?
While our initial analysis in §3.3 suggests that any type of data could work based on the strong correlation between VLM performance and class occurrence in LLaVA-training data, which is generally VQA-type rather than classification-focused, we further investigate how data types affect final performance in this section.
In previous experiments detailed in §3.2, we fine-tuned the VLM using a classification-focused template (e.g., “<image> Question: What type of object is in this photo? Answer: <class name>.”). Here, we fine-tune the VLM on captioning data. This captioning data is generated by providing an image and the ground-truth class name to GPT4 and requesting it to generate a caption that includes the ground-truth name. Specifically, the prompt for GPT4 is: “Generate a short caption for the image. The caption must include the ground-truth label of the image, which is <class name>.” The generated captions are diverse for each class. For example, for the “guinea pig” class, captions could be “A guinea pig playing in the grass.” or “A close-up view of a guinea pig in an indoor environment.” We fine-tune the VLM on this caption-focused data using the same experimental settings and compare the model performance with that fine-tuned on classification-focused data. Success is defined as whether the generated caption for a validation image contains the ground-truth class name.
The performances of fine-tuned LLaVA1.5-7B on three datasets are reported in Table 10. Our findings indicate that data is the main determining factor for VLM performance, and the data type does not matter much. From Table 10, we can see that fine-tuning on both classification-focused data and caption-focused data achieves a similar level of accuracy on Flowers102, StanfordCars, and Caltech101, at least 48.6% higher than the non-fine-tuned model. This indicates that to enable a VLM to perform classification on a specific class, it is unnecessary to use classification-specific data. Instead, any data related to that class can be used, provided that the data includes the class name.
Model Training Data Type [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
LLaVA1.5-7B Zero-shot - 5.9 0.0 47.1
CLIP-L Fine-tuned Classification Data 98.6 91.5 97.6
LLaVA1.5-7B Fine-tuned Classification Data 97.6 90.4 97.5
LLaVA1.5-7B Fine-tuned Captioning Data 92.0 85.4 95.7
Table 10: Analysis of data types. We fine-tune the VLM on the caption-focused data generated by GPT4 using the same experimental settings and find that data is the main determining factor for VLM performance, and the data type does not matter much.

### v2

@@ Abstract
Image classification is one of the most fundamental capabilities of machine vision intelligence. In this work, we revisit the image classification task using visually-grounded language models (VLMs) such as GPT-4V and LLaVA. We find that existing proprietary and public VLMs, despite often using CLIP as a vision encoder and having many more parameters, significantly underperform CLIP on standard image classification benchmarks like ImageNet. To understand the reason, we explore several hypotheses concerning the inference algorithms, training objectives, and data processing in VLMs. Our analysis reveals that the primary cause is data-related: critical information for image classification is encoded in the VLM’s latent space but can only be effectively decoded with enough training data. Specifically, there is a strong correlation between the frequency of class exposure during VLM training and instruction-tuning and the VLM’s performance in those classes; when trained with sufficient data, VLMs can match the accuracy of state-of-the-art classification models. Based on these findings, we enhance a VLM by integrating classification-focused datasets into its training, and demonstrate that the enhanced classification performance of the VLM transfers to its general capabilities, resulting in an improvement of 11.8% on the newly collected ImageWikiQA dataset. [^1] Project page: https://yuhui-zh15.github.io/VLMClassifier-Website/.
@@ 1 Introduction
In summary, our contributions are threefold:
  • Evaluating VLM classification weaknesses: Our evaluations using ten VLMs across four benchmarks show that VLMs significantly lag behind CLIP in classification, uncovering a gap unaddressed by previous research.
  • Analyzing causes of poor classification performance: Testing various hypotheses, we find that the lack of alignment data, rather than information lost or training objective, is the primary reason for VLMs’ underperformance in classification.
  • Enhancing VLMs with classification data: By adding classification data, we improve VLMs’ accuracy in object recognition and overall performance, supporting that classification capability is foundational for complex object-related reasoning.
@@ 2 VLMs are Bad at Image Classification
2.1 Models
VLMs.
CLIPs.
2.2 Data
2.3 Evaluation Protocol
VLMs.
CLIPs.
2.4 Results
@@ 3 Why are VLMs Bad Image Classifiers?
    Finally, we investigate data-related questions. For example, does the VLM training data include enough classification data and cover enough classes? We find a strong correlation between class exposure in training and model performance. Moreover, VLMs can achieve the same level of performance as CLIPs when trained with enough data. These results suggest that data is the primary cause and effective solution for the poor classification performance of VLMs.
@@ 3.1 Inference
Table 2: Analysis of VLMs from the inference perspective. (Top) We explore prompt variation such as wording, label order, chain-of-thought and find it has limited impact on the performance. (Bottom) We leverage the probabilistic inference strategy, which improves the performance but still fails to close the gap between VLMs and CLIPs. Results are from the official validation set.
@@ 3.2 Training
Table 3: Analysis of VLMs from the training perspective. (Left) We conduct feature probing experiments on the VLM’s last layer and find that the information required for classification is mostly preserved in the VLM’s latent space. (Right) We fine-tune VLMs on the classification datasets using the text generation objective and find that the text generation training objective is as effective as the traditional cross-entropy for learning classification, which eliminates the VLM-CLIP performance gap, with VLMs now being the state-of-the-art classifier. Results are from the official validation set.
@@
We find that the text generation training objective is as effective as the traditional cross-entropy for learning classification tasks, which eliminates the performance gap between VLMs and CLIPs, with VLMs now being the state-of-the-art classifier. From Table 3, we find that LLaVA1.5-7B achieves 85.7% accuracy on ImageNet, same as 85.2% accuracy when fine-tuning its vision encoder CLIP-L. The same findings apply to all the VLMs on all the datasets.
@@ 3.3 Data
As observed, the visual information is preserved in the VLM’s latent space, and the VLM’s training objective is sufficient for learning classification tasks. We hypothesize that the poor classification performance of VLMs is due to their training data. For example, there might be insufficient classification data or a lack of diverse classes. To investigate this, we analyze the LLaVA1.5-7B training data [32], the only fully publicly available VLM training dataset. We examined the relationship between the frequency of class occurrence and the VLM’s classification performance on those classes [^2] LLaVA training contains two stages. Here we combine the data from the two stages. More in Appendix B.6..
Our findings indicate that data determines VLM classification performance. As shown in Figure 3, there is a strong correlation between the presence of class labels in the VLM training data and the VLM’s classification accuracy for those classes. The LLaVA1.5-7B model achieves 82.7% zero-shot accuracy on ImageNet classes with more than 10,000 occurrences in its training data, but only 3.0% accuracy on classes with fewer than 10 occurrences. The Spearman correlation between class frequency and class performance is very high (0.76 on ImageNet; see Appendix §B.6). In contrast, there is no correlation between its vision encoder CLIP-L’s performance on those classes with the same VLM training data, and after fine-tuning LLaVA1.5-7B on the ImageNet classification data, the strong correlation disappears. Combining all of our results, we conclude that data plays a critical role in determining the VLM’s classification performance.
@@
Given that data availability is critical for VLM classification performance, we now examine whether the data type specifically impacts results. For instance, to train a VLM to classify a class like “guinea pig,” should it rely on classification-focused data (e.g., “<image> What type of object is in this photo? Guinea pig”) or can it use broader types, such as caption-focused data (e.g., “<image> Generate a caption for the image. A guinea pig playing in the grass”) or VQA-focused data (e.g., “<image> What is the native region of this guinea pig? South America”)?
In previous experiments (see §3.2), we fine-tuned the VLM with a classification-focused template. Here, we fine-tune the VLM on caption-focused data generated by prompting GPT4 to produce captions incorporating each class’s ground-truth name. Specifically, we used the prompt: “Generate a short caption for the image. The caption must include the ground-truth label of the image, which is <class name>.” This approach produces diverse captions per class, such as “A guinea pig playing in the grass” or “A close-up view of a guinea pig in an indoor environment.” We then fine-tune the VLM on this caption-focused data using identical experimental settings and evaluate it against the model trained with classification-focused data. Success is determined by whether the generated caption for a validation image includes the ground-truth class name.
We find that data presence is the primary factor for VLM classification performance, with data type playing a minimal role. From Table 4, we observe that models fine-tuned on both classification- and caption-focused data perform similarly on Flowers102, StanfordCars, and Caltech101, achieving accuracy gains of at least 48.6% over the non-fine-tuned model. This suggests that enabling a VLM to classify a specific class does not require classification-specific data; any data containing the class name suffices.
These findings emphasize a critical data-centric view of VLM training. A common question for VLM training is whether alignment between vision encoder outputs and text generation necessitates multi-modal data with specific class labels, despite both vision and language components seeing these classes in uni-modal pre-training. Previous studies have suggested that this alignment stage may be data-efficient or even unnecessary [32, 18], likely due to their evaluation settings not requiring vision-centric fine-grained recognition. Our work, however, shows that 1) multi-modal data is essential for alignment and 2) increasing data quantity linearly boosts performance. Concurrent work from DeepMind supports our findings, showing that most VLM tasks benefit from pre-training on larger datasets [6].
Model Fine-tuning Data Type [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
LLaVA1.5-7B Zero-shot - 5.9 0.0 47.1
CLIP-L Fine-tuned Classification Data 98.6 91.5 97.6
LLaVA1.5-7B Fine-tuned Classification Data 97.6 90.4 97.5
LLaVA1.5-7B Fine-tuned Captioning Data 92.0 85.4 95.7
Table 4: Analysis of data types. We fine-tune the VLM on the caption-focused data generated by GPT4 using the same experimental settings as Table 3 and find that data is the main determining factor for VLM performance, and the data type does not matter much.
@@ 4.2 ImageWikiQA
To bridge the gap between classification and advanced capabilities, we introduce ImageWikiQA, an object-centric, knowledge-intensive question answering dataset that combines both worlds. Each question in ImageWikiQA is a multiple-choice question with four options and one correct answer (e.g., “<image showing a guinea pig> What is the native region of this object? A. South America; B. Africa; C. Asia; D. Australia”). Although each question does not directly ask for the category of the object within the image, the question can only be accurately answered if the class of the object is correctly identified. Example questions from ImageWikiQA can be found in Figure 1 (right) and Appendix §C.1.
@@ Acknowledgements
Appendix A Supplementary for Section 2
@@ A.2 Data
A.3 Synonymous Text Labels in ImageNet Evaluation
In ImageNet, many classes have multiple valid textual labels. For instance, the class identifier n01496331 corresponds to both “electric ray” and “crampfish”. To address this, we incorporated label synonyms [^3] https://gist.github.com/yrevar/942d3a0ac09ec9e5eb3a in our evaluation. We treat the classification as correct if any textual label is included in the VLM generation. Including synonyms improves accuracy by only 1%-3% (Table 8), still leaving a notable performance gap when compared to CLIP.
Model Accuracy w/o Synonyms Accuracy w/ Synonyms Improvement ($\Delta$)
BLIP2-2.7B 25.3 27.8 2.5
IBLIP-7B 14.6 16.5 1.9
IBLIP-13B 14.7 16.6 1.9
LLaVA1.5-7B 22.8 24.6 1.8
LLaVANeXT-V7B 29.4 32.2 2.8
LLaVA1.5-13B 24.3 26.0 1.7
LLaVANeXT-M7B 32.3 35.1 2.8
Claude3 53.6 56.3 2.7
GeminiPro 39.2 42.5 3.3
GPT4 48.5 51.1 2.6
Table 8: Model accuracy with and without label synonyms on ImageNet, and corresponding improvement.
@@ B.1 Prompt Variation Detail
For the fixed order, we use the default dataset label order for each image.
@@ B.2 Label Set Size Details
The table version of Figure 2 is provided in Table 9. By reducing the number of labels for classification, we can narrow the gap between VLM and CLIP, but a gap remains across all label set sizes, even with just two labels (two-way classification).
@@
We find that while the absolute gap between VLMs and CLIPs narrows with reduced label size, the relative gap increases. For example, in two-way classification on ImageNet, VLMs have a 5.7% error rate, while CLIP has a 0.2% error rate, resulting in a 28.5x gap; for 20 classes, VLMs have an 18.0% error rate, while CLIP has a 2.5% error rate, resulting in a 7.2x gap. This trend is indicated by the error rates between VLMs and CLIP shown in Figure 4.
Larger gap on fine-grained classification datasets.
The absolute gap on Flowers and Cars is larger than ImageNet and Caltech when reducing the label set size. This may be because Flowers and Cars are more fine-grained classification datasets, while ImageNet and Caltech are more coarse-grained. VLMs are weaker in fine-grained classification compared to CLIP, resulting in a larger gap.
@@ B.6 Data Detail
B.7 Analysis of Proprietary VLMs
Method LLaVA1.5-7B GPT4
Prompt Variation
Base Prompt 22.8 5.9 0.0 47.1 48.5 51.0 0.1 61.0
w/ Label (Fixed Order) N/A 6.1 0.1 70.5 61.2 79.0 57.1 91.2
w/ Label (Random Order) N/A 10.2 0.0 62.1 60.6 79.9 58.2 94.2
w/ Label (Random Order) + CoT N/A 18.1 0.4 64.5 62.2 79.0 59.8 94.0
Table 12: Analysis of VLMs from the inference perspective. Compared to Table 2, we added the results for the proprietary GPT4 model. We explore prompt variation such as label order and chain-of-thought (CoT) and find it has a limited impact on the performance, which applies to proprietary VLMs such as GPT4.
Figure 6: Analysis of the label set size. Compared to Figure 2, we added the line for the proprietary GPT4 model. LLaVA1.5-7B uses CLIP-L as its vision encoder, while GPT4 presumably employs a superior vision encoder. Therefore, comparing GPT4 to CLIP-L is not entirely fair. However, we observe that GPT4 performs worse than CLIP-L on ImageNet even when the number of classes is reduced to 2.
In the main paper, we analyzed two public VLMs about why they are bad at image classification. For proprietary VLMs, they haven’t released the training details and data usage, so it is impossible to directly address training-related or data-related hypotheses.
For inference-related hypotheses, we conduct additional experiments with GPT4 (Table 12, Figure 6), and the conclusion is the same as the public VLMs. For example, GPT-4 achieves 60.6% accuracy on ImageNet without CoT. With CoT, GPT-4 achieves 62.2% accuracy, demonstrating that CoT has a limited impact on image classification, even for large VLMs.
For training-related and data-related hypotheses, recent work PaliGemma [6] from DeepMind shows that most VLM tasks benefit significantly from longer pre-training with more data (in its Figure 4 and Appendix K). This aligns with our main conclusion that data is the critical factor in improving VLM performance.
B.8 Open-world vs. Closed-world Setting
§3 and §4 all use “open-world” settings (not providing the label set in the prompt), except for prompt variation analysis and label set size analysis in §3.1, using “closed-world” settings (providing the label set in the prompt).
Naturally, the “closed-world” setting narrows the model’s generation space and increases accuracy. When the model can easily predict the class in the label set by modifying the inference space (probabilistic inference in §3.1 and probing VLM in §3.2) or training VLMs with more classification data (§4), the advantage of the “closed-world” setting doesn’t exist. Thus, for these experiments, we use the “open-world” setting.
B.9 Number Comparison of Table 1 vs. Table 2
Table 2’s “Base Prompt w/ Label (Random Order)” is equivalent to Table 1’s “Closed-World Setting”. In this setting, we concatenate label candidates in a random order for each question: “What’s in the image? Choose one from random_shuffle([cat, dog, pig])”. The accuracy for these two settings is identical between Table 2 and Table  1.
Table 2’s “Base Prompt w/ Label (Fixed Order)” concatenates label candidates in a fixed order for each question: “What’s in the image? Choose one from [cat, dog, pig]”. This setting is used to rule out the possibility that the order of labels might affect model performance.
B.10 Why Do VLMs Have Information but Cannot Decode?
In 3, we find that VLMs have the essential information for classification. If the VLMs have the information, then why are they not able to decode it?
One possible reason is that the VLM’s decoding space is too large and not aligned with the visual features. Specifically, VLM decoding is performed through next-word prediction, which usually involves a vocabulary of over 10,000 words. The output text embedding is not aligned with the visual features due to insufficient data to align these spaces.
In general, having information in a model does not necessarily mean the model can express that information. This phenomenon is also observed in other research areas. For example, after training ResNets or ViTs with self-supervised learning methods like SimCLR, the model acquires discriminative information for different classes. However, this information can only be decoded by adding and training a linear layer on a new dataset.
Similarly, we demonstrate that VLMs possess classification information, but it is not readily expressible. Adding classification-related data helps bridge the gap between “information possession” and “information expression” (refer to Table 3).
B.11 Why Does Fine-tuning Only on ImageNet Get Worse Performance on ImageWikiQA?
In Table 5, the “Finetuned on ImageNet” model performance (30.6) is worse than the non-finetuned model performance (38.0). This is because fine-tuning often improves in-distribution performance but does not necessarily enhance out-of-distribution or general capabilities.
Table 5 presents the results of models fine-tuned on ImageNet and evaluated on ImageWikiQA, which are vastly different datasets. ImageNet questions only require classification, such as “Q: <image> What is in the image? A: dog,” while ImageWikiQA questions demand both classification and knowledge/reasoning, such as “Q: <image> What is the native region of this object? A: South America.”
Fine-tuning solely on ImageNet trains the model to classify, but it can lead to a loss of general capabilities like reasoning and knowledge, resulting in lower performance on ImageWikiQA (30.6 fine-tuned vs. 38.0 pre-trained). This phenomenon is known as “catastrophic forgetting” [19].
However, when we fine-tune on both ImageNet and instruction tuning data, the model learns classification while retaining its original capabilities, leading to a significant performance improvement on ImageWikiQA (49.8 fine-tuned vs. 38.0 pre-trained).
B.12 Analysis of Linear Probed VLM from Data Perspective
Figure 7 plots results for the linear probed LLaVA model, which aligns closely with the fine-tuned LLaVA. This alignment shows a zero correlation with class frequency since fine-tuning or probing on the balanced ImageNet dataset equalizes class frequencies, where each class is evenly represented.
Figure 7: Analysis of linear probed VLM from the data perspective. Compared to Figure 3, we added the line for the linear probed LLaVA model, which has zero correlation. The zero correlation is expected because probing on the ImageNet dataset equalizes class frequencies, as ImageNet classes are evenly distributed.
@@ C.1 ImageWikiQA Details
Data collection details.
ImageWikiQA only contains a test set, providing a valuable benchmark for assessing both fine-grained classification and knowledge-based reasoning in VLMs.
The ImageWikiQA dataset was constructed in two stages. In the first stage, we generated questions at the class level for each of the 1,000 ImageNet classes. For instance, questions like “What is the native region of <class name>?” were created to capture high-level class information. In the second stage, we refined these questions at the image level. For each ImageNet class, we randomly sampled 3 images from the 50 available images from the ImageNet validation set and formulated specific questions like “What is the native region of <image i>?”. Each question in ImageWikiQA is paired with a single image, ensuring only one image per question.
The ImageWikiQA dataset contains 2,000 unique questions. A question like “What is the native region of <image i>?” is counted multiple times if it is asked for different images or classes. For example, if the question is asked for 2 cat images and 2 dog images, it would count as 4 questions in total. To avoid over-representing any particular question, a question like “What is the native region of <image i>?” is limited to a maximum of 3 occurrences across the 2,000 questions.
Data generation prompt.
@@ C.2 Results Details
C.3 Performance of Fine-tuned VLM on Other Datasets
To understand whether incorporating the ImageNet 1.28M classification data into the original LLaVA instruction-tuning data will harm the model’s general capability, we further evaluated the performance of our fine-tuned visually-grounded language model (VLM) on additional benchmarks to ensure its robustness across a variety of tasks. In particular, we tested the model on TextVQA [42], POPE [29], and MMVet [51]. The results, presented in Table 13, show that the fine-tuned LLaVA1.5-7B maintains almost identical performance compared to the original version.
This result aligns with expectations since our fine-tuning process incorporated all of the original LLaVA training data during instruction tuning. For instance, on TextVQA, the fine-tuned model achieved an accuracy of 58.0%, closely matching the original model’s 58.2%. Additionally, the fine-tuned model slightly outperformed the original on POPE Popular and POPE Adverse benchmarks, showing small improvements of 0.2% and 0.3%, respectively. Performance on MMVet remained unchanged at 31.1%. These findings demonstrate that the fine-tuning process preserves the model’s strengths across these varied benchmarks.
Model TextVQA POPE Popular POPE Adverse MMVet
LLaVA1.5-7B (Official Released) 58.2 86.1 84.2 31.1
LLaVA1.5-7B (Further Finetuned) 58.0 86.3 84.5 31.1
Table 13: Performance of LLaVA1.5-7B before and after fine-tuning on TextVQA, POPE, and MMVet datasets. Fine-tuning resulted in consistent performance across all benchmarks.
C.4 Low Accuracy on ImageWikiQA with Ground-truth Class Names
Table 5 presents the accuracy on ImageWikiQA, which includes questions derived from Wikipedia requiring extensive world knowledge (e.g., “What is the native region of the guinea pig?”).
The modest accuracy of 55.9% by LLaVA1.5-7B with provided ground-truth class names reflects its limited world knowledge for ImageNet classes. This limitation aligns with findings that smaller models, like Vicuna-7B, lack world knowledge compared to larger models like GPT-4.
C.5 Other Fine-Tuning and Inference Strategies
Other fine-tuning or inference modifications, such as applying a linear probing loss in the VLM output space during training or using k-nearest neighbors (KNN) in the VLM output space for evaluation, can also enhance VLM classification performance.
However, our primary goal is to improve VLMs’ general capabilities across a range of tasks, not just classification. The universal inference interface for various tasks is text generation, whereas approaches like KNN-based inference are tailored solely to classification and do not generalize to other task types. For fine-tuning, we experimented with adding a new token and applying a linear probing loss. This approach did not yield improvements on ImageWikiQA, indicating that task-specific fine-tuning does not enhance the overall capabilities of VLMs.
In summary, without changing VLM training or inference to maintain its general capabilities for different tasks, adding data is the most promising and effective approach.
Appendix D Summary of Contributions
Object recognition is fundamental to the general capability of visually-grounded language models (VLMs), yet current VLMs perform poorly in this area. We are the first to thoroughly investigate this critical issue and propose a potential solution to improve it.
Our primary contribution lies in identifying the problem (i.e., VLMs are inadequate image classifiers, a significant weakness that has been overlooked) and conducting an in-depth investigation (i.e., understanding why VLMs are poor image classifiers and how to address this issue).
In summary, our contributions are threefold:
  • Thorough Evaluation to Identify the Problem: We thoroughly evaluated current public and proprietary VLMs on four common classification benchmarks and discovered that their performance significantly lags behind CLIP. This finding is counterintuitive because VLMs often use CLIP as their vision encoders. This finding reveals a weakness in VLMs that previous works have not widely noted [2, 49]. Understanding the limitations of VLMs is crucial given their increasing deployment in various scenarios.
  • Hypothesis Testing to Understand the Problem: Given the poor performance of VLMs in classification, we investigated the underlying reasons. We considered multiple plausible hypotheses related to VLM inference, training, and data perspectives. For example, essential information for classification could be lost during the vision encoder’s propagation through multiple LLM layers; VLMs might be inherently poor at classification due to their text generation training objective compared to the standard cross-entropy objective. Our thorough investigation ruled out these reasons but revealed the major issue: the lack of alignment data. This finding is counterintuitive because other alternative hypotheses also seemed plausible.
  • Improving VLMs based on Our Understanding: Based on our findings, we explored ways to enhance VLM performance. We believe that classification is foundational to more advanced capabilities. For example, if a VLM cannot accurately classify mushroom species, it will also struggle with follow-up questions, such as whether a particular mushroom is poisonous. Indeed, we found that simply adding classification data not only improves VLMs’ classification performance but also their general capabilities, demonstrating that accurately identifying objects is a prerequisite for answering complex questions about these objects.
Impact:
Recent work from Google DeepMind, PaliGemma [6], supports our main conclusion that data is the critical factor in improving VLM performance. They also found that most VLM tasks benefit significantly from longer pre-training with more data (Appendix K). This demonstrates that our analysis can inspire researchers to build better VLMs in the future.